OpenStudy (anonymous):

What the hell Complete the following ordered pairs for the equation. 4x + 5y = -22 (_, -2) (2, _) (7, _)

OpenStudy (anonymous):

(_, -2) (x, y) they give you the y-value: -2 so when y = -2 4x + 5y = -22 4x + 5(-2) = -22 4x - 10 = -22 +10 +10 ----- ---- 4x = -12 --- ---- 4 4 x = ?

OpenStudy (anonymous):

that will give you "x" for the first ordered pair (_, -2)

OpenStudy (anonymous):

@Moonraven what did you get?

OpenStudy (anonymous):

yea